Assets, Beginning of Period 12,000

Liabilities, Beginning of Period 5,000

Assets, End of Period 23,000

Liabilities, End of Period 9,000

Investment by Owner 10,000

Revenues 99,000

Withdrawals by Owner 0

Solve for the expenses incurred during the period.
1
Expert's answer
2020-01-21T07:36:11-0500

Costs (costs, expenses) are a decrease in assets or an increase in liabilities (liabilities) or a combination of both in connection with the production and delivery of goods and services in the framework of the enterprise. The expenses associated with the implementation of capital and financial investments do not relate to the expenses of the enterprise.

Therefore, we calculate the expenses:

Expenses = (Assets, End of Period - Assets, Beginning of Period) +(Liabilities End of Period - Liabilities, Beginning of Period)=(23000-12000)+(9000-5000)=11000+4000=15000